Subject: 3.5-m Users Committee minutes 11/03/08
From: Bruce Gillespie
Submitted: Mon, 10 Nov 2008 22:34:28 -0700
Message number: 1117
(previous: 1116,
next: 1118
up: Index)
APO 3.5-m Users Committee Phonecon, 11/03/08
Attending: Suzanne Hawley, Bruce Gillespie, Scott Anderson, Al
Harper, Meredith Drosback (for Remy Indebetouw), Mark Klaene, Bill
Ketzeback, John Bally, Rene Walterbos, and Michael Strauss
Absent: Jon Holtzman, Jon Fulbright
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
Princeton (Michael Strauss) - Michael said that the PU users have been
happy with their early experiences using TripleSpec.
Colorado (John Bally) - John is presently at APO hosting a CU student
field trip. Regarding NIC-FPS, he suggested that we should
standardize the instrument's focus script by limiting Fowler sampling
to 0 or 1 (to save time), and that bright stars should be used for
focussing with a narrow-band filter (to minimize persistent image
problems inherent to the detector). The persistent images, which
appear as negative relative signal, take a long time and many flushes
to remove them. Al Harper reported that he has recently looked at NIC-
FPS images, and sees latent images as positive, too--he will send an
example to John Bally. Suzanne asked Mark to discuss the suggestion
of standardizing NIC-FPS focussing procedures with the Observing
Specialists.
Chicago (Al Harper) - Al had no further comments on user feedback.
NMSU (Rene Walterbos) - Rene reported that the spectrograph focus of
DIS is recently noticeably worse than it has been earlier in the year;
this is particularly detrimental to radial velocity measurements.
Several weeks ago, Russet McMillan and Jon Holtzman recently adjusted
the DIS focus, as we routinely do seasonally. Mark said that Russet
then rechecked the DIS focus yesterday, and found that it needed a
significant readjustment that improved the resolution by a factor of
~2. This is surprising, and may be related to the recent spate of
warm weather we have been having. Suzanne said that DIS
spectrographic focus should be properly maintained for all users, not
just for programs requiring the best resolution. She asked Mark to
look into the feasibility of making a monthly check and adjustment of
DIS spectrographic focus, rather than quarterly.
Washington (Scott Anderson) - Scott mentioned that Agile users are
reporting frequent network disconnections, which interrupts the data
reads in high-speed observing scripts. During a disconnection, the
software repeatedly tries to reconnect, which to the APO system looks
like a hacking attack. The APO software then denies the requests from
the remote user, which exacerbates the situation. As a temporary fix,
Fritz Stauffer has adjusted the parameters for access denials, but a
better solution will come when Agile is run under TUI; TUI will allow
that data collection to continue during network disconnections.
Virginia (Meredith Drosback) - nothing to report.
Johns Hopkins (Jon Fulbright) - no report.
**********************************
Discussion of telescope/instruments report - Klaene:
Mark summarized the report, noting the recent good weather and
relatively trouble-free telescope and instrument performance. There
have been minor bugs in the software controlling the NA2 instrument
rotator, and some of the motion-control parameters for the tertiary
mirror have been tweaked. The cranky DIS-blue ion pump is being
watched, and we may choose to replace it during an upcoming spell of
bad weather. The recent work and repairs to TripleSpec have been
completed, and the Agile move to the TR2 port is underway, although we
have had some problems with motor torque with the new image rotator.
The APOLLO lunar-ranging program has been running well, and the plan
for next year's CIF projects is being finalized.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 10/2/08
through 10/28/08
Mark Klaene
0) Overview
Under mostly clear night skies, the telescope has been working quite
well during the period. Minor NA2 rotator problems have caused small
losses in observing time. The new boiler for lower dorm is complete
and the new site-wide fire alarm system is receiving its final
adjustments.
1) Telescope
New power supplies were installed to address the axial stepper motor
noise and seemed to have helped, although occasional on-full-step-
error's are still being seen. The instrument rotator for the TR2 port
is being installed this week, and tuning will be done on its servo
loop once it is installed. Axis controller software continues to be
refined on the NA2 rotator axis.
We saw our first freeze for this season, which caused a problem with
the DIS slitviewer shutter. The system has since been winterized by
installing insulation around the camera. This shutter has shown cold
temperature problems in previous years and replacement units do not
seem to help, while the installation of thermal insulation seemed to
fix the problem last year.
2) Instruments
The DIS-blue ion pump is occasionally running high and the dewar may
need to be opened up in the future for servicing--spare ion pumps are
now on site. DIS-red and SPIcam are fully operational with no problems.
Echelle and NIC-FPS both continue to be operational with no problems.
We are still waiting for UVa to design and build up a new NIC-FPS fan-
out board that will hopefully address the residual readout noise.
TripleSpec warmup, gauge replacement and ion pump installation are now
complete. The instrument is cold and fully operational.
Agile is still available at NA2; it will be moved to the TR2 port once
the new rotator is fully tested. TUI integration for Agile
continues. Agile spares procurement and documentation updates are
also in progress.
GFP is currently warm because it is not scheduled for observing until
late November.
APOLLO remains operational with more record-breaking amounts of lunar-
return photons. A high-precision gravimeter (needed by APOLLO to
determine the exact position of the telescope) is slated for arrival
late in November, and plans are in progress to install it at the 3.5-m.
3) CIF projects
Planning for next year's CIF projects is well underway. The principal
projects are direct drives for Az and altitude axes, new axis
controllers for az and alt, and cal lamp upgrades.
**********************************
Calibration (Flat, Arc lamps) upgrades - group:
Suzanne said that as part of the CIF plan for next year, we are
considering various improvements to the on-telescope calibration lamps
system. We would like to hear suggestions from the users. Rene asked
if one upgrade would be to get an arc lamp with more lines in the
blue. Suzanne said yes, and also for a flat-field source with more
flux in the blue and ultraviolet. Scott said that he asked and got no
replies from the UW users. In any case, Suzanne said that we will do
something that improves the line density and flat-field flux in the
blue, now that we have improved DIS sensitivity in that wavelength
region. She asked if anyone knew if better calibration sources were
needed for NIC-FPS. John Bally said that NIC-FPS users generally use
OH emission in the night sky for calibration. In closing the topic,
Suzanne asked that users give us their suggestions for improving the
calibration lamps system; this opportunity will only stay open over
the next few months.
**********************************
Autofocus status - Klaene:
Mark pointed out that autofocus has been a project worked on this past
year. Using the NA2 guider to sense changes in focus, the system has
been prototyped and works reasonably well, but work is still needed to
characterize and improve it. Users can ask the Observing Specialists
to turn the system on during their observing programs, and feedback is
welcome. Suzanne added that the principal use of the autofocus system
is to keep targets in focus during DIS observations--we will need to
implement autofocus differently for TripleSpec and SPIcam. One
consequence of the system is that it is OK to guide on field stars
using autofocus. We will soon have user information on autofocus on
our web pages.
**********************************
Communication with users during observing - Hawley:
Suzanne remarked that some users find the TUI message window
inadequate for enabling easy communication with the Observing
Specialists during remote observing runs. Over the years, several
Observing Specialists have opposed the introduction of a live A/V link
to the control room, but it seems reasonable to consider the use of
live audio, such as iChat or Skype, to facilitate user-ObsSpec voice
communication. Bill pointed out that the link might need to be muted
most of the time, to avoid a dueling of the music played at the
respective ends of the hookup. The message window does have the
advantage of being logged, but it could still be used for purposes
requiring an archive record, and for instances of cut-and-paste
communications. Suzanne said that the site will look into
implementing a trial version of live audio, probably Skype, between
users and the 3.5-m's control room.
**********************************
Future instrumentation (priorities for replacements, upgrades) - Hawley:
Suzanne announced that we will have an opportunity beginning in 2010
to provide some financial support for the development of new and
upgraded instrumentation for the 3.5-m telescope. When the 2009 CIF
projects are completed, instrumentation and their upgrades will be the
higher priority for CIF funds. Specifically, both the echelle and
SPIcam are probably nearing the end of their useful lives in their
current configuration, and both instruments have seen dwindling user
demand in the past few years. In order to get the next-generation
instruments on the telescope, we need both user involvement and
leadership--if someone in the ARC user community wants to step forward
and take on the PI role of defining and developing new instrumentation
for the telescope, funding can be made available. Suzanne will take
this proposal to the ARC Board of Governors in two weeks, and she
would be interested in hearing any user feedback both before and after
the Board meeting (slh at astro.washington.edu).
**********************************
3.5-m publications list update - Walterbos:
Rene compiles the list of new publications and theses based on 3.5-m
data annually for the ARC Board meeting. He asks that the Users
Committee members remind their users to send him listings of all
publication and theses from the period Oct. 07 through Oct. 08 (rwalterb@nmsu.edu
).
**********************************
ACTION ITEMS:
[open from previous months]:
ACTION: Committee members should send one or two science highlights
from the telescope users from their institutions to Suzanne (slh at
astro.washington.edu) and Bruce (gillespi at apo.nmsu.edu) not later
than 1 October for use in the Annual Report to the ARC Board of
Governors meeting in November.
STATUS: Completed.
[new actions from last month]:
ACTION: The Users Committee representatives are to ask their
respective users for suggestions on improving the calibration lamp
system on the telescope, e.g., flux, wavelength distribution, emission-
line calibration, etc., and report at the next meeting.
STATUS: Discussed with Users Committee 11/3.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday, 8
December, at 8:30 AM Pacific Time. The agenda and other materials
will be sent to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1117 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO